OpenStudy (anonymous):

After this problem is simplified, what will the denominator be? [x^2/3]^2

OpenStudy (anonymous):

@iGreen

OpenStudy (muzzack):

Final result : \[\frac{ x ^{4} }{3^{4}}\]

OpenStudy (igreen):

Is it this? \(\Large\frac{x^2}{3}^2\)

OpenStudy (anonymous):

muzzack are you sure?

OpenStudy (muzzack):

sorry \[\frac{ x^4 }{3^2 }\]

OpenStudy (igreen):

Yeah, which simplifies to: \(\Large\frac{x^4}{9}\)

OpenStudy (anonymous):

\[\left[ \frac{ x ^{2} }{ 3 } \right]^{2}\]

OpenStudy (anonymous):

That's the equation.

OpenStudy (anonymous):

Took me a while to get that down, still learning how to work this.

OpenStudy (anonymous):

So the answer is 9?

OpenStudy (igreen):

Yes.